# Generalized hydrodynamics of integrable quantum circuits

### v1

@@ B.1 The boundary quantum transfer matrix construction
$|\Psi_{0}(z)\rangle=\left(\frac{z^{1/2}|01\rangle-z^{-1/2}|10\rangle}{\sqrt{|z|+|z|^{-1}}}\right)^{\otimes L/2}$ (88)
@@
The BQTM approach allows to compute the GGE rapidity distributions from initial states of the form (88) by recasting the discrete time evolution in terms of auxiliary transfer matrices acting in the space direction, and where the initial state enters as a boundary condition.
@@
$\xi_{\pm}=\arctan\left(\frac{z+1}{z-1}\tan\left(\tfrac{x}{2}\pm\tfrac{\gamma}{2}\right)\right)$ (89)
(as oppposed to the case of most generic two-site product states considered in [86], the states (88) correspond to diagonal boundary conditions, so all the other parameters vanish).
@@
$\mathcal{T}(u)=\omega_{1}(u)+\omega_{2}(u)\,,$ (90)
@@ B.2 The T–system and Y–system
$\mathcal{T}_{j}$ $=$ $\mathcal{T}_{j-1}^{-}\mathcal{T}_{1}^{[j-1]}-f^{[j-3]}\mathcal{T}_{j-2}^{[-2]}\,,\qquad j\geq 2,$ (92)
where for any function $F(u)$ we have introduced the short-hand notation $F^{[k]}(u)\equiv F(u+ik\gamma/2)$, and where $f(u-\gamma/2)=\omega_{1}(u+\gamma/2)\omega_{2}(u-\gamma/2)$. In the more general case where the transfer matrix acts on a chain of $N$ spins, or for more general (e.g. non-diagonal) boundary conditions, the definition of $f$ is changed, but the general structure (92) remains the same [86].
@@
$\mathcal{T}_{j}^{[m]}\mathcal{T}_{j}^{[-m]}$ $=\mathcal{T}_{j+m}\mathcal{T}_{j-m}$
$+$ $\Phi_{j-m+1}\mathcal{T}_{m-1}^{[j+1]}\mathcal{T}_{m-1}^{[-j-1]}\,,$ (93)
@@
$Y_{j}(\lambda)=\frac{\mathcal{T}_{j-1}(u)\mathcal{T}_{j+1}(u)}{\Phi_{j}(u)}\,,\qquad j\geq 1\,.$ (94)
@@
$Y_{j}\left(u+\frac{\gamma}{2}\right)Y_{j}\left(u-\frac{\gamma}{2}\right)=\left[1+Y_{j+1}\left(u\right)\right]\left[1+Y_{j-1}\left(u\right)\right]\,,$ (95)
@@
$1+Y_{1}(u)=\left(1+\mathfrak{a}(u-\gamma/2)\right)\left(1+1/\mathfrak{a}(u+\gamma/2)\right)\,,$ (96)
@@
$\mathfrak{a}(u)$ $=\frac{\omega_{1}(u)}{\omega_{2}(u)}$
$=$ $\frac{\sin(2u+\gamma)}{\sin(2u-\gamma)}\frac{\sin(u+\xi_{+}-\gamma/2)\sin(u+\xi_{-}-\gamma/2)}{\sin(u-\xi_{+}+\gamma/2)\sin(u-\xi_{-}+\gamma/2)}\,.$ (97)
@@ B.3 Truncation of the T– and Y– system at $\gamma=\frac{\pi}{p+1}$
$\mathcal{T}_{p+1}(u)=\alpha(u)\mathcal{T}_{p-1}(u)+\beta(u)\,,$ (98)
@@
$\alpha(u)$ $=$ $\frac{\left(\Omega^{[-p-2]}\Omega_{1}^{[-p-2]}\Omega_{2}^{[-p-2]}\right)^{2}}{\Phi_{p}}\,,$ (99)
$\beta(u)$ $=$ $\Omega^{-[p+2]}\left((\Omega_{1}^{[-p-2]})^{2}+(\Omega_{2}^{[-p-2]})^{2}\right)\,,$ (100)
@@
As a result of (98), the Y–system can also be truncated: introducing the function
$K(u)=\frac{\Omega^{[-p-2]}\Omega_{1}^{[-p-2]}\Omega_{2}^{[-p-2]}}{\Phi_{p}}\mathcal{T}_{p-1}(u)\,,$ (102)
@@
$\begin{split}Y_{j}^{+}Y_{j}^{-}&=(1+Y_{j+1})(1+Y_{j-1})\,,\quad j=1,\ldots p-1\,,\\ 1+Y_{p-1}&=K^{+}K^{-}\,,\\ 1+Y_{p}&=1+(\mathfrak{b}+\mathfrak{b}^{-1})K+K^{2}\,,\qquad\mathfrak{b}=\frac{\Omega_{1}^{[-p-2]}}{\Omega_{2}^{[-p-2]}}\,.\end{split}$ (103)
B.4 The GGE densities at $\gamma=\tfrac{\pi}{p+1}$
@@
$\begin{split}\eta_{j}(\lambda)&=Y_{j}(i\lambda)\,,\qquad j=1,\ldots p-1\\ \eta_{p}(\lambda)&=\mathfrak{b}(i\lambda)K(i\lambda)\\ \eta_{p+1}(\lambda)&=\mathfrak{b}(i\lambda)/K(i\lambda)\end{split}$ (104)
@@ Appendix C Details on the free-fermion points of the model
From Eqs. (74) and (75), we see that the unitary operators $U_{e}(\tau)$, $U_{o}(\tau)$ are written as the exponential of quadratic Hamiltonians. This makes it possible to diagonalise the Floquet operator $U(\tau)$ by means of an appropriate Bogoliubov rotation [33]. In addition, the mapping to free fermions makes it possible to simulate efficiently the dynamics of the system, using the formalism of Gaussian states [96]. We used this method to produce the numerical data presented in Sec. IV.1, as we explain in the following.
@@
$\Gamma_{ij}=\langle\psi|c^{\dagger}_{i}c_{j}|\psi\rangle\,.$ (105)
@@
$\Gamma(2t+1)$ $=V_{o}^{\ast}\Gamma(2t)V_{o}^{T}\,,$ (106)
$\Gamma(2t)$ $=V_{e}^{\ast}\Gamma(2t-1)V_{e}^{T}\,,$ (107)
@@
$(A_{o})_{2n,2n+1}=(A_{o})_{2n+1,2n}=-i\frac{\tau}{2}\,,$ (109)
$(A_{e})_{2n-1,2n}=(A_{e})_{2n,2n-1}=-i\frac{\tau}{2}\,.$ (110)
We can also map the initial states (41)–(43) into fermionic Fock states, which are Gaussian. This is true also for bipartite initial states, where the two halves of the system are prepared in different product states chosen among  (41)–(43). In fact, it is straightforward to compute their covariance matrix. Given the initial matrix $\Gamma$, the evolution is computed numerically via iteration of Eqs. (106).

### v2

@@ B.1 The boundary quantum transfer matrix construction
$|\Psi_{0}(z)\rangle=\left(\frac{z^{1/2}|01\rangle-z^{-1/2}|10\rangle}{\sqrt{|z|+|z|^{-1}}}\right)^{\otimes L/2}$ (B.1)
@@
The BQTM approach allows to compute the GGE rapidity distributions from initial states of the form (B.1) by recasting the discrete time evolution in terms of auxiliary transfer matrices acting in the space direction, and where the initial state enters as a boundary condition.
@@
$\xi_{\pm}=\arctan\left(\frac{z+1}{z-1}\tan\left(\tfrac{x}{2}\pm\tfrac{\gamma}{2}\right)\right)$ (B.2)
(as oppposed to the case of most generic two-site product states considered in [87], the states (B.1) correspond to diagonal boundary conditions, so all the other parameters vanish).
@@
$\mathcal{T}(u)=\omega_{1}(u)+\omega_{2}(u)\,,$ (B.3)
@@ B.2 The T–system and Y–system
$\mathcal{T}_{j}$ $=$ $\mathcal{T}_{j-1}^{-}\mathcal{T}_{1}^{[j-1]}-f^{[j-3]}\mathcal{T}_{j-2}^{[-2]}\,,\qquad j\geq 2,$ (B.5)
where for any function $F(u)$ we have introduced the short-hand notation $F^{[k]}(u)\equiv F(u+ik\gamma/2)$, and where $f(u-\gamma/2)=\omega_{1}(u+\gamma/2)\omega_{2}(u-\gamma/2)$. In the more general case where the transfer matrix acts on a chain of $N$ spins, or for more general (e.g. non-diagonal) boundary conditions, the definition of $f$ is changed, but the general structure (B.5) remains the same [87].
@@
$\mathcal{T}_{j}^{[m]}\mathcal{T}_{j}^{[-m]}$ $=\mathcal{T}_{j+m}\mathcal{T}_{j-m}+\Phi_{j-m+1}\mathcal{T}_{m-1}^{[j+1]}\mathcal{T}_{m-1}^{[-j-1]}\,,$ (B.6)
@@
$Y_{j}(\lambda)=\frac{\mathcal{T}_{j-1}(u)\mathcal{T}_{j+1}(u)}{\Phi_{j}(u)}\,,\qquad j\geq 1\,.$ (B.7)
@@
$Y_{j}\left(u+\frac{\gamma}{2}\right)Y_{j}\left(u-\frac{\gamma}{2}\right)=\left[1+Y_{j+1}\left(u\right)\right]\left[1+Y_{j-1}\left(u\right)\right]\,,$ (B.8)
@@
$1+Y_{1}(u)=\left(1+\mathfrak{a}(u-\gamma/2)\right)\left(1+1/\mathfrak{a}(u+\gamma/2)\right)\,,$ (B.9)
@@
$\mathfrak{a}(u)$ $=\frac{\omega_{1}(u)}{\omega_{2}(u)}=\frac{\sin(2u+\gamma)}{\sin(2u-\gamma)}\frac{\sin(u+\xi_{+}-\gamma/2)\sin(u+\xi_{-}-\gamma/2)}{\sin(u-\xi_{+}+\gamma/2)\sin(u-\xi_{-}+\gamma/2)}\,.$ (B.10)
@@ B.3 Truncation of the T– and Y– system at $\gamma=\frac{\pi}{p+1}$
$\mathcal{T}_{p+1}(u)=\alpha(u)\mathcal{T}_{p-1}(u)+\beta(u)\,,$ (B.11)
@@
$\alpha(u)$ $=$ $\frac{\left(\Omega^{[-p-2]}\Omega_{1}^{[-p-2]}\Omega_{2}^{[-p-2]}\right)^{2}}{\Phi_{p}}\,,$ (B.12)
$\beta(u)$ $=$ $\Omega^{-[p+2]}\left((\Omega_{1}^{[-p-2]})^{2}+(\Omega_{2}^{[-p-2]})^{2}\right)\,,$ (B.13)
@@
As a result of (B.11), the Y–system can also be truncated: introducing the function
$K(u)=\frac{\Omega^{[-p-2]}\Omega_{1}^{[-p-2]}\Omega_{2}^{[-p-2]}}{\Phi_{p}}\mathcal{T}_{p-1}(u)\,,$ (B.15)
@@
$\begin{split}Y_{j}^{+}Y_{j}^{-}&=(1+Y_{j+1})(1+Y_{j-1})\,,\quad j=1,\ldots p-1\,,\\ 1+Y_{p-1}&=K^{+}K^{-}\,,\\ 1+Y_{p}&=1+(\mathfrak{b}+\mathfrak{b}^{-1})K+K^{2}\,,\qquad\mathfrak{b}=\frac{\Omega_{1}^{[-p-2]}}{\Omega_{2}^{[-p-2]}}\,.\end{split}$ (B.16)
B.4 The GGE densities at $\gamma=\pi/(p+1)$
@@
$\begin{split}\eta_{j}(\lambda)&=Y_{j}(i\lambda)\,,\qquad j=1,\ldots p-1\\ \eta_{p}(\lambda)&=\mathfrak{b}(i\lambda)K(i\lambda)\\ \eta_{p+1}(\lambda)&=\mathfrak{b}(i\lambda)/K(i\lambda)\end{split}$ (B.17)
@@ Appendix C Details on the free-fermion points of the model
From Eqs. (74) and (75), we see that the unitary operators $U_{e}(\tau)$, $U_{o}(\tau)$ are written as the exponential of quadratic Hamiltonians. This makes it possible to diagonalise the Floquet operator $U(\tau)$ by means of an appropriate Bogoliubov rotation [33]. In addition, the mapping to free fermions makes it possible to simulate efficiently the dynamics of the system, using the formalism of Gaussian states [97]. We used this method to produce the numerical data presented in Sec. 4.1, as we explain in the following.
@@
$\Gamma_{ij}=\langle\psi|c^{\dagger}_{i}c_{j}|\psi\rangle\,.$ (C.1)
@@
$\Gamma(2t+1)$ $=V_{o}^{\ast}\Gamma(2t)V_{o}^{T}\,,$ (C.2)
$\Gamma(2t)$ $=V_{e}^{\ast}\Gamma(2t-1)V_{e}^{T}\,,$ (C.3)
@@
$(A_{o})_{2n,2n+1}=(A_{o})_{2n+1,2n}=-i\frac{\tau}{2}\,,$ (C.5)
$(A_{e})_{2n-1,2n}=(A_{e})_{2n,2n-1}=-i\frac{\tau}{2}\,.$ (C.6)
We can also map the initial states (41)–(43) into fermionic Fock states, which are Gaussian. This is true also for bipartite initial states, where the two halves of the system are prepared in different product states chosen among  (41)–(43). In fact, it is straightforward to compute their covariance matrix. Given the initial matrix $\Gamma$, the evolution is computed numerically via iteration of Eqs. (C.2).